Let A be a square matrix, and AT denotes the transpose of A. Show that the following hold true.
(a) (AT)T=A
(b) (A+B)T=AT+BT
(c) (AB)T=BT AT.
1
Expert's answer
2020-12-01T06:08:32-0500

Let "A=\left[
\begin{array}{cccc}
a_{11} & a_{12} & \ldots & a_{1n}\\
a_{21} & a_{22} & \ldots & a_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1} & a_{n2} & \ldots & a_{nn}
\end{array}
\right]" and "B=\left[
\begin{array}{cccc}
b_{11} & b_{12} & \ldots & b_{1n}\\
b_{21} & b_{22} & \ldots & b_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
b_{n1} & b_{n2} & \ldots & b_{nn}
\end{array}
\right]".



(a) "A^T=\left[
\begin{array}{cccc}
a_{11} & a_{21} & \ldots & a_{n1}\\
a_{12} & a_{22} & \ldots & a_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
a_{1n} & a_{2n} & \ldots & a_{nn}
\end{array}
\right]" and "(A^T)^T=\left[
\begin{array}{cccc}
a_{11} & a_{12} & \ldots & a_{1n}\\
a_{21} & a_{22} & \ldots & a_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1} & a_{n2} & \ldots & a_{nn}
\end{array}
\right]=A."



(b) "(A+B)^T=\left(\left[
\begin{array}{cccc}
a_{11} & a_{12} & \ldots & a_{1n}\\
a_{21} & a_{22} & \ldots & a_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1} & a_{n2} & \ldots & a_{nn}
\end{array}
\right]+

\left[
\begin{array}{cccc}
b_{11} & b_{12} & \ldots & b_{1n}\\
b_{21} & b_{22} & \ldots & b_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
b_{n1} & b_{n2} & \ldots & b_{nn}
\end{array}
\right]\right)^T="


"=\left[
\begin{array}{cccc}
a_{11}+b_{11} & a_{12}+b_{12} & \ldots & a_{1n}+b_{1n}\\
a_{21}+b_{21} & a_{22}+b_{22} & \ldots & a_{2n}+b_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1}+b_{n1} & a_{n2}+b_{n2} & \ldots & a_{nn}+b_{nn}
\end{array}
\right]^T="


"=
\left[
\begin{array}{cccc}
a_{11}+b_{11} & a_{21}+b_{21} & \ldots & a_{n1}+b_{n1}\\
a_{12}+b_{12} & a_{22}+b_{22} & \ldots & a_{n2}+b_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
a_{1n}+b_{1n} & a_{2n}+b_{2n} & \ldots & a_{nn}+b_{nn}
\end{array}
\right]="


"=\left[
\begin{array}{cccc}
a_{11} & a_{21} & \ldots & a_{n1}\\
a_{12} & a_{22} & \ldots & a_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
a_{1n} & a_{2n} & \ldots & a_{nn}
\end{array}
\right]+

\left[
\begin{array}{cccc}
b_{11} & b_{21} & \ldots & b_{n1}\\
b_{12} & b_{22} & \ldots & b_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
b_{1n} & b_{2n} & \ldots & b_{nn}
\end{array}
\right]=A^T+B^T."


(c) "(AB)^T=\left(\left[
\begin{array}{cccc}
a_{11} & a_{12} & \ldots & a_{1n}\\
a_{21} & a_{22} & \ldots & a_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1} & a_{n2} & \ldots & a_{nn}
\end{array}
\right]\cdot

\left[
\begin{array}{cccc}
b_{11} & b_{12} & \ldots & b_{1n}\\
b_{21} & b_{22} & \ldots & b_{2n}\\
\ldots & \ldots & \ldots & \ldots\\
b_{n1} & b_{n2} & \ldots & b_{nn}
\end{array}
\right]\right)^T="


"=\left[
\begin{array}{cccc}
a_{11}b_{11}+\cdots+a_{1n}b_{n1} & a_{11}b_{12}+\cdots+a_{1n}b_{n2} & \ldots & a_{11}b_{1n}+\cdots+a_{1n}b_{nn}\\
a_{21}b_{11}+\cdots+a_{2n}b_{n1} & a_{21}b_{12}+\cdots+a_{2n}b_{n2} & \ldots & a_{21}b_{1n}+\cdots+a_{2n}b_{nn}\\
\ldots & \ldots & \ldots & \ldots\\
a_{n1}b_{11}+\cdots+a_{nn}b_{n1} & a_{n1}b_{12}+\cdots+a_{nn}b_{n2} & \ldots & a_{n1}b_{1n}+\cdots+a_{nn}b_{nn}\\
\end{array}
\right]^T="


"=\left[
\begin{array}{cccc}
a_{11}b_{11}+\cdots+a_{1n}b_{n1} & a_{21}b_{11}+\cdots+a_{2n}b_{n1} & \ldots & a_{n1}b_{11}+\cdots+a_{nn}b_{n1}\\
a_{11}b_{12}+\cdots+a_{1n}b_{n2} & a_{21}b_{12}+\cdots+a_{2n}b_{n2} & \ldots & a_{n1}b_{12}+\cdots+a_{nn}b_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
a_{11}b_{1n}+\cdots+a_{1n}b_{nn} & a_{21}b_{1n}+\cdots+a_{2n}b_{nn} & \ldots & a_{n1}b_{1n}+\cdots+a_{nn}b_{nn}\\
\end{array}
\right]="


"=
\left[
\begin{array}{cccc}
b_{11} & b_{21} & \ldots & b_{n1}\\
b_{12} & b_{22} & \ldots & b_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
b_{1n} & b_{2n} & \ldots & b_{nn}
\end{array}
\right]
\cdot
\left[
\begin{array}{cccc}
a_{11} & a_{21} & \ldots & a_{n1}\\
a_{12} & a_{22} & \ldots & a_{n2}\\
\ldots & \ldots & \ldots & \ldots\\
a_{1n} & a_{2n} & \ldots & a_{nn}
\end{array}
\right]
=B^T A^T."